# Tight Bounds for Schrödinger Potential Estimation in Unpaired Data Translation

### v2

@@ Appendix D Log-potential variance and conditional variance bounds
Appendix E Orlicz norm bounds
In this section, we present upper bounds on the $\psi_{1}$-norm of the log-potential difference $\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})$ as well as of $\log\big({\mathcal{T}}_{T}[e^{\varphi}](X_{0})/{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})\big)$. This is a starting point in our analysis of the empirical risk minimizer $\widehat{\varphi}$ defined in (11). These results play a crucial role in the proof of Lemma F.1 about uniform concentration of the empirical means
@@ Appendix E Orlicz norm bounds
around their expectation as they allow us to use Bernstein’s inequality for sub-exponential random variables [Lecué and Mitchell, 2012, Proposition 5.2].
Lemma E.1.
@@
The proofs of Lemmata E.1 and E.2 are deferred to Appendices E.1 and E.2, respectively. We also provide an upper bound on the Orlicz norm of $\log\big(\pi^{*}(X_{T}^{*}\,|\,x_{0})/\pi^{\varphi}(X_{T}^{*}\,|\,x_{0})\big)$ with respect to the conditional density $\pi^{*}(\cdot\,|\,x_{0}))$. Unlike Lemmata E.1 and E.2, we use Lemma E.3 in verification of a Bernstein-type condition for the empirical KL-loss. To be more precise, we combine this result with Lemma 4.1 to obtain an upper bound on the expected conditional variance $\mathbb{E}\mathrm{Var}\left(\varphi^{*}(X_{T}^{*})-\varphi(X_{T}^{*})\,\big|\,X_{0}\right)$ (see Lemma D.1 for the details).
Lemma E.3.
@@
The proof of Lemma E.3 is postponed to Appendix E.3.
E.1 Proof of Lemma E.1
@@
$\left\|\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\right\|_{\psi_{1}}\leqslant\left\|L\left\|\Sigma^{-1/2}(X_{T}^{*}-m)\right\|^{2}+M\right\|_{\psi_{1}}\leqslant L\left\|\left\|\Sigma^{-1/2}(X_{T}^{*}-m)\right\|^{2}\right\|_{\psi_{1}}+\frac{M}{\log 2}.$
In the last inequality, we used the fact that $\|c\|_{\psi_{1}}\leqslant|c|/\log 2$ for any constant $c\in\mathbb{R}$. The rest of the proof is devoted to the study of the Orlicz norm of $\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}$. According to Lemma F.3, for any $x_{0}\in\mathrm{supp}(\rho_{0})$, it holds that
$\mathbb{E}\left[e^{b\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}=x_{0}\right]\leqslant 2^{d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)+2be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
@@
$\mathbb{E}e^{b\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}$ $=\mathbb{E}\mathbb{E}\left[e^{b\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}\right]$
$\leqslant\frac{2^{d/2}}{\big(1-e^{-2bT}\big)^{10e^{2}\sqrt{d}}}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{K}}(T)\left(\frac{be^{2}R^{2}}{\sqrt{d}}+4e^{2}\sqrt{d}\right)\arcsin(e^{-bT})+2be^{-2bT}R^{2}\right\}.$
@@
E.2 Proof of Lemma E.2
According to Lemma F.10, for any $x\in\mathbb{R}^{d}$, it holds that
@@
$\mathbb{E}\left[\exp\left\{\lambda\left|\log\frac{\pi^{*}(X_{T}^{*}\,|\,x_{0})}{\pi^{\varphi}(X_{T}^{*}\,|\,x_{0})}\right|\right\}\,\Big|\,X_{0}=x_{0}\right]\leqslant\sqrt{1+\mathbb{E}\left[\exp\left\{2\lambda\log\frac{\pi^{*}(X_{T}^{*}\,|\,x_{0})}{\pi^{\varphi}(X_{T}^{*}\,|\,x_{0})}\right\}\,\Big|\,X_{0}=x_{0}\right]}.$
@@
$\leqslant\exp\left\{\frac{d\log 2}{2}+M\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big(2{\mathcal{K}}(T)+1\big)+2be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
@@
Moreover, according to Lemma F.10, we have
@@
Summing up (26) and (E.3), we obtain that
@@
$\mathbb{E}\left[\exp\left\{2\lambda\log\frac{\pi^{*}(X_{T}^{*}\,|\,x_{0})}{\pi^{\varphi}(X_{T}^{*}\,|\,x_{0})}\right\}\,\Big|\,X_{0}=x_{0}\right]\leqslant e^{6\lambda M+2\lambda{\mathcal{A}}(x_{0},T)({\mathcal{K}}(T)+1)}\;\mathbb{E}\left[e^{2\lambda L\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}=x_{0}\right].$
Applying Lemma F.3 and taking into account that $\lambda=\min\{1/2,b/(2L)\}\leqslant b/(2L)$, we conclude that
@@
$\leqslant\exp\left\{\frac{d\log 2}{2}+M\big({\mathcal{K}}(T)+6\lambda\big)+{\mathcal{A}}(x_{0},T)\big({\mathcal{K}}(T)+2\lambda{\mathcal{K}}(T)+2\lambda\big)+4e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}$
$\leqslant\exp\left\{\frac{d\log 2}{2}+M\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big(2{\mathcal{K}}(T)+1\big)+2be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
@@
$\leqslant 2^{(d+1)/2}\exp\left\{\frac{M}{2}\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big({\mathcal{K}}(T)+1/2\big)+be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
This inequality yields the desired upper bound on the Orlicz norm of $\log\big(\pi^{*}(X_{T}^{*}\,|\,x_{0})/\pi^{\varphi}(X_{T}^{*}\,|\,x_{0})\big)$. Indeed, let us take
$v=\frac{\log 2}{\lambda}\vee\frac{\log 2}{\lambda}\left(\frac{d+1}{2}+\frac{M}{2}\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big({\mathcal{K}}(T)+1/2\big)+be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right).$
@@
$\leqslant\left(2^{(d+1)/2}\exp\left\{\frac{M}{2}\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big({\mathcal{K}}(T)+1/2\big)+be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}\right)^{1/(\lambda v)}\leqslant 2.$
@@
Appendix F Auxiliary results
This section collects auxiliary results. We provide their proofs (except for Lemma F.10) in Appendices F.1F.9. For the proof of Lemma F.10, a reader is referred to [Puchkin et al., 2025, Lemma B.3].
Lemma F.1.
@@ Appendix F Auxiliary results
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$ $\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
@@
$\left|\frac{1}{n}\sum\limits_{j=1}^{n}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}-\mathbb{E}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\right|$ $\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
@@
$\leqslant e^{{\mathcal{A}}(x,T)({\mathcal{K}}(T)+1/{\mathcal{K}}(T))+M{\mathcal{K}}(T)}\left(M+\frac{LR^{2}}{2}+\frac{L(50d+d^{2}/8)}{b}\right)^{1-1/{\mathcal{K}}(T)}\left({\mathcal{T}}_{\infty}|f_{1}-f_{0}|\right)^{1/{\mathcal{K}}(T)}.$
Lemma F.3.
@@
$\mathbb{E}\left[e^{2\lambda L\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}=x_{0}\right]\leqslant 2^{d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)+4e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
Lemma F.4.
@@
$\log\mathbb{E}\left[e^{u^{\top}\Sigma^{-1/2}(X_{T}^{*}-m)}\,\big|\,X_{0}=x_{0}\right]\leqslant M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)+e^{-bT}\,u^{\top}\Sigma^{-1/2}(x_{0}-m)+\frac{(1-e^{-2bT})\|u\|^{2}}{4b},$
@@
are sub-exponential. Second, we apply [Lecué and Mitchell, 2012, Proposition 5.2] to quantify large deviation bounds simultaneously for all $\varphi\in\Phi_{\varepsilon}$. Finally, we extend these bounds to concentration inequalities holding uniformly for all $\varphi\in\Phi$.
@@
They follow from Assumption 2, 3, and 4. Indeed, according to Lemma E.1, under Assumptions 3 and 4, it holds that
$\left\|\varphi^{*}(Y_{i})-\varphi(Y_{i})\right\|_{\psi_{1}}\lesssim\left(1\vee\frac{L}{b}\right)\left(M+d+e^{-bT}\left(\frac{R^{2}}{\sqrt{d}}+\sqrt{d}+bR^{2}\right)\right)\lesssim\left(1\vee\frac{L}{b}\right)(M+d)\quad\text{for all $1\leqslant i\leqslant n$.}$ (32)
On the other hand, Lemma E.2 implies that
$\left\|\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}\right\|_{\psi_{1}}\leqslant\left(\frac{be^{2}R^{2}\arcsin(e^{-bT})}{\sqrt{d}\log 2}+\frac{2\log{\mathcal{K}}(T)}{\log 2}\right)\left(\frac{1}{{\mathcal{K}}(T)}+{\mathcal{K}}(T)\right)+\frac{M{\mathcal{K}}(T)}{\log 2}\quad\text{for all $1\leqslant j\leqslant n$,}$
where ${\mathcal{K}}(T)$ is defined in (31). Let us note that there exists $T_{0}\lesssim(\log d\vee\log R)/b$ such that for all $T\geqslant T_{0}$
@@
Step 2: concentration of sub-exponential random variables.  The $\psi_{1}$-norm bounds obtained on the previous step allow us to use Bernstein’s inequality for unbounded random variables. Let us fix an arbitrary $\varphi\in\Phi_{\varepsilon}$ and apply [Lecué and Mitchell, 2012, Proposition 5.2]). Then, for any $\delta\in(0,1)$, with probability at least $1-\delta/(4|\Phi_{\varepsilon}|)$, it holds that
$\quad+\left\|\max\limits_{1\leqslant i\leqslant n}\big(\varphi^{*}(Y_{i})-\varphi(Y_{i})\big)\right\|_{\psi_{1}}\cdot\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}.$ (34)
@@
$\left|\frac{1}{n}\sum\limits_{j=1}^{n}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}-\mathbb{E}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\right|$ $\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$
$\quad+\left\|\max\limits_{1\leqslant j\leqslant n}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}\right\|_{\psi_{1}}\cdot\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}$ (35)
Due to the the union bound, there exists an event ${\mathcal{E}}_{0}$ of probability at least $1-\delta/2$ such that the inequalities (F.1) and (F.1) hold simultaneously for all $\varphi\in\Phi_{\varepsilon}$ on this event. Moreover, using Pisier’s inequality (see, for example, [Lecué and Mitchell, 2012, p. 1827]), we conclude that
@@
$\left|\frac{1}{n}\sum\limits_{j=1}^{n}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}-\mathbb{E}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\right|$ $\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$
@@
on the event ${\mathcal{E}}_{0}$ simultaneously for all $\varphi\in\Phi_{\varepsilon}$. In view of (32) and (33), we have
$\quad+\left(1\vee\frac{L}{b}\right)\frac{(M+d)\log(4|\Phi_{\varepsilon}|/\delta)\log n}{n}.$ (36)
@@
$\left|\frac{1}{n}\sum\limits_{j=1}^{n}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}-\mathbb{E}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\right|$ $\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$
$\quad+\frac{(M+1)\log(4|\Phi_{\varepsilon}|/\delta)\log n}{n}$ (37)
@@
Step 3: from $\varepsilon$-nets to uniform bounds.  The goal of this step is to transform this upper bound to a one holding uniformly for all $\varphi\in\Phi$. According to [Rigollet and Hütter, 2023, Theorem 1.19], the norm of a sub-Gaussian random vector satisfies the inequality

### v3

@@ Appendix C Log-potential variance and conditional variance bounds
Appendix D Orlicz norm bounds
In this section, we present upper bounds on the $\psi_{1}$-norm of the log-potential difference $\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})$ as well as of $\log\big({\mathcal{T}}_{T}[e^{\varphi}](X_{0})/{\mathcal{T}}_{T}[e^{\varphi^{*}}](X_{0})\big)$. This is a starting point in our analysis of the empirical risk minimizer $\widehat{\varphi}$ defined in (12). These results play a crucial role in the proof of Lemma E.1 about uniform concentration of the empirical means
@@ Appendix D Orlicz norm bounds
around their expectation as they allow us to use Bernstein’s inequality for sub-exponential random variables (Lecué & Mitchell, 2012, Proposition 5.2).
Lemma D.1.
@@
The proofs of Lemmata D.1 and D.2 are deferred to Appendices D.1 and D.2, respectively. We also provide an upper bound on the Orlicz norm of $\log\big(\pi^{*}(X_{T}^{*}\,|\,x_{0})/\pi^{\varphi}(X_{T}^{*}\,|\,x_{0})\big)$ with respect to the conditional density $\pi^{*}(\cdot\,|\,x_{0}))$. Unlike Lemmata D.1 and D.2, we use Lemma D.3 in verification of a Bernstein-type condition for the empirical KL-loss. To be more precise, we combine this result with Lemma B.1 to obtain an upper bound on the expected conditional variance $\mathbb{E}\mathrm{Var}\left(\varphi^{*}(X_{T}^{*})-\varphi(X_{T}^{*})\,\big|\,X_{0}\right)$ (see Lemma C.1 for the details).
Lemma D.3.
@@
The proof of Lemma D.3 is postponed to Appendix D.3.
D.1 Proof of Lemma D.1
@@
$\left\|\varphi(X_{T}^{*})-\varphi^{*}(X_{T}^{*})\right\|_{\psi_{1}}$ $\leqslant\left\|L\left\|\Sigma^{-1/2}(X_{T}^{*}-m)\right\|^{2}+M\right\|_{\psi_{1}}$
$\leqslant L\left\|\left\|\Sigma^{-1/2}(X_{T}^{*}-m)\right\|^{2}\right\|_{\psi_{1}}+\frac{M}{\log 2}.$
In the last inequality, we used the fact that $\|c\|_{\psi_{1}}\leqslant|c|/\log 2$ for any constant $c\in\mathbb{R}$. The rest of the proof is devoted to the study of the Orlicz norm of $\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}$. According to Lemma E.3, for any $x_{0}\in\mathrm{supp}(\rho_{0})$, it holds that
$\mathbb{E}\left[e^{b\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}=x_{0}\right]$ $\leqslant 2^{d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)\right\}$
$\quad\cdot\exp\left\{2be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
@@
$\mathbb{E}e^{b\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}$
$=\mathbb{E}\mathbb{E}\left[e^{b\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}\right]$
$\leqslant\frac{2^{d/2}}{\big(1-e^{-2bT}\big)^{10e^{2}\sqrt{d}}}$
$\quad\cdot\exp\left\{M{\mathcal{K}}(T)+{\mathcal{K}}(T)\left(\frac{be^{2}R^{2}}{\sqrt{d}}+4e^{2}\sqrt{d}\right)\arcsin(e^{-bT})+2be^{-2bT}R^{2}\right\}.$
@@
D.2 Proof of Lemma D.2
According to Lemma E.10, for any $x\in\mathbb{R}^{d}$, it holds that
@@
$\leqslant\sqrt{1+\mathbb{E}\left[\exp\left\{2\lambda\log\frac{\pi^{*}(X_{T}^{*}\,|\,x_{0})}{\pi^{\varphi}(X_{T}^{*}\,|\,x_{0})}\right\}\,\Big|\,X_{0}=x_{0}\right]}.$
@@
$\leqslant\exp\left\{\frac{d\log 2}{2}+M\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big(2{\mathcal{K}}(T)+1\big)\right\}$
$\quad\cdot\exp\left\{2be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
@@
Moreover, according to Lemma E.10, we have
@@
Summing up (34) and (D.3), we obtain that
@@
$\leqslant e^{6\lambda M+2\lambda{\mathcal{A}}(x_{0},T)({\mathcal{K}}(T)+1)}\;\mathbb{E}\left[e^{2\lambda L\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}=x_{0}\right].$
Applying Lemma E.3 and taking into account that $\lambda=\min\{1/2,b/(2L)\}\leqslant b/(2L)$, we conclude that
$\leqslant\exp\left\{\frac{d\log 2}{2}+M\big({\mathcal{K}}(T)+6\lambda\big)+{\mathcal{A}}(x_{0},T)\big({\mathcal{K}}(T)+2\lambda{\mathcal{K}}(T)+2\lambda\big)\right\}$
$\quad\cdot\exp\left\{4e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}$
$\leqslant\exp\left\{\frac{d\log 2}{2}+M\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big(2{\mathcal{K}}(T)+1\big)\right\}$
$\quad\cdot\exp\left\{2be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
@@
$\leqslant 2^{(d+1)/2}\exp\left\{\frac{M}{2}\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big({\mathcal{K}}(T)+1/2\big)\right\}$
$\quad\cdot\exp\left\{be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
This inequality yields the desired upper bound on the Orlicz norm of $\log\big(\pi^{*}(X_{T}^{*}\,|\,x_{0})/\pi^{\varphi}(X_{T}^{*}\,|\,x_{0})\big)$. Indeed, let $v$ be the maximum of $\log(2)/\lambda$ and
$\frac{\log 2}{\lambda}\left(\frac{d+1}{2}+\frac{M}{2}\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big({\mathcal{K}}(T)+1/2\big)+be^{-2bT}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right).$
@@
$\leqslant\left(2^{(d+1)/2}\exp\left\{\frac{M}{2}\big({\mathcal{K}}(T)+3\big)+{\mathcal{A}}(x_{0},T)\big({\mathcal{K}}(T)+1/2\big)\right\}\right)^{1/(\lambda v)}$
$\quad\cdot\exp\left\{\frac{be^{-2bT}}{\lambda v}\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}$
$\leqslant 2.$
@@
Appendix E Auxiliary results
This section collects auxiliary results. We provide their proofs (except for Lemma E.10) in Appendices E.1E.9. For the proof of Lemma E.10, a reader is referred to (Puchkin et al., 2025, Lemma B.3).
Lemma E.1.
@@ Appendix E Auxiliary results
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\left(\varphi(Y_{i})-\varphi^{*}(Y_{i})\right)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$
$\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
@@
$\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\big(D\log(\Lambda nd)+\log(1/\delta)\big)}{n}}$
@@
$\leqslant e^{{\mathcal{A}}(x,T)({\mathcal{K}}(T)+1/{\mathcal{K}}(T))+M{\mathcal{K}}(T)}$
$\quad\cdot\left(M+\frac{LR^{2}}{2}+\frac{L(50d+d^{2}/8)}{b}\right)^{1-1/{\mathcal{K}}(T)}\left({\mathcal{T}}_{\infty}|f_{1}-f_{0}|\right)^{1/{\mathcal{K}}(T)}.$
Lemma E.3.
@@
$\mathbb{E}\left[e^{2\lambda L\|\Sigma^{-1/2}(X_{T}^{*}-m)\|^{2}}\,\Big|\,X_{0}=x_{0}\right]$ $\leqslant 2^{d/2}\exp\left\{M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)\right\}$
$\quad\exp\left\{4e^{-2bT}\lambda L\left\|\Sigma^{-1/2}(x_{0}-m)\right\|^{2}\right\}.$
Lemma E.4.
@@
$\log\mathbb{E}\left[e^{u^{\top}\Sigma^{-1/2}(X_{T}^{*}-m)}\,\big|\,X_{0}=x_{0}\right]$
$\leqslant M{\mathcal{K}}(T)+{\mathcal{A}}(x_{0},T){\mathcal{K}}(T)+e^{-bT}\,u^{\top}\Sigma^{-1/2}(x_{0}-m)+\frac{(1-e^{-2bT})\|u\|^{2}}{4b},$
@@
are sub-exponential. Second, we apply (Lecué & Mitchell, 2012, Proposition 5.2) to quantify large deviation bounds simultaneously for all $\varphi\in\Phi_{\varepsilon}$. Finally, we extend these bounds to concentration inequalities holding uniformly for all $\varphi\in\Phi$.
@@
They follow from Assumption 2, 3, and 4. Indeed, according to Lemma D.1, under Assumptions 3 and 4, it holds that
$\left\|\varphi^{*}(Y_{i})-\varphi(Y_{i})\right\|_{\psi_{1}}$ $\lesssim\left(1\vee\frac{L}{b}\right)\left(M+d+e^{-bT}\left(\frac{R^{2}}{\sqrt{d}}+\sqrt{d}+bR^{2}\right)\right)$
$\lesssim\left(1\vee\frac{L}{b}\right)(M+d)\quad\text{for all $1\leqslant i\leqslant n$.}$ (40)
On the other hand, Lemma D.2 implies that
$\left\|\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}\right\|_{\psi_{1}}\leqslant\left(\frac{be^{2}R^{2}\arcsin(e^{-bT})}{\sqrt{d}\log 2}+\frac{2\log{\mathcal{K}}(T)}{\log 2}\right)\left(\frac{1}{{\mathcal{K}}(T)}+{\mathcal{K}}(T)\right)+\frac{M{\mathcal{K}}(T)}{\log 2}$
for all $1\leqslant j\leqslant n$, where ${\mathcal{K}}(T)$ is defined in (39). Let us note that there exists $T_{0}\lesssim(\log d\vee\log R)/b$ such that for all $T\geqslant T_{0}$
@@
Step 2: concentration of sub-exponential random variables.  The $\psi_{1}$-norm bounds obtained on the previous step allow us to use Bernstein’s inequality for unbounded random variables. Let us fix an arbitrary $\varphi\in\Phi_{\varepsilon}$ and apply (Lecué & Mitchell, 2012, Proposition 5.2)). Then, for any $\delta\in(0,1)$, with probability at least $1-\delta/(4|\Phi_{\varepsilon}|)$, it holds that
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\big(\varphi^{*}(Y_{i})-\varphi(Y_{i})\big)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$
$\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$ (42)
$\quad+\left\|\max\limits_{1\leqslant i\leqslant n}\big(\varphi^{*}(Y_{i})-\varphi(Y_{i})\big)\right\|_{\psi_{1}}\cdot\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}.$
@@
$\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$ (43)
$\quad+\left\|\max\limits_{1\leqslant j\leqslant n}\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{j})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{j})}\right\|_{\psi_{1}}\cdot\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}.$
Due to the the union bound, there exists an event ${\mathcal{E}}_{0}$ of probability at least $1-\delta/2$ such that the inequalities (E.1) and (E.1) hold simultaneously for all $\varphi\in\Phi_{\varepsilon}$ on this event. Moreover, using Pisier’s inequality (see, for example, (Lecué & Mitchell, 2012, p. 1827)), we conclude that
@@
$\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$
@@
on the event ${\mathcal{E}}_{0}$ simultaneously for all $\varphi\in\Phi_{\varepsilon}$. In view of (E.1) and (41), we have
$\left|\frac{1}{n}\sum\limits_{i=1}^{n}\big(\varphi^{*}(Y_{i})-\varphi(Y_{i})\big)-\mathbb{E}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\right|$
$\lesssim\sqrt{\frac{\mathrm{Var}\big(\varphi^{*}(Y_{1})-\varphi(Y_{1})\big)\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$ (44)
$\quad+\left(1\vee\frac{L}{b}\right)\frac{(M+d)\log(4|\Phi_{\varepsilon}|/\delta)\log n}{n}.$
@@
$\lesssim\sqrt{\mathrm{Var}\left(\log\frac{{\mathcal{T}}_{T}[e^{\varphi}](Z_{1})}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](Z_{1})}\right)\frac{\log(4|\Phi_{\varepsilon}|/\delta)}{n}}$ (45)
$\quad+\frac{(M+1)\log(4|\Phi_{\varepsilon}|/\delta)\log n}{n}$
@@
Step 3: from $\varepsilon$-nets to uniform bounds.  The goal of this step is to transform this upper bound to a one holding uniformly for all $\varphi\in\Phi$. According to (Rigollet & Hütter, 2023, Theorem 1.19), the norm of a sub-Gaussian random vector satisfies the inequality